OpenStudy (anonymous):

I'm working through an example and I get stuck here: 1/(x+sqrt{x^2 +1}) (1 + (2x /(2sqrt{x^2 +1})) it ends as 1/(sqrt{x^2 +1}) I don't know how to manipulate the second part of the first expression. Thanks for your insight.

OpenStudy (phi):

A little hard to read: \[\frac{1}{x+\sqrt{x^2+1}}\left( 1+\frac{2x}{2\sqrt{x^2+1}} \right)\] cancel the 2's for the first fraction, multiply top and bottom by x-sqrt(x^2+1) this makes the denominator x^2 - (x^2+1) = 1 put the 1 in the 2nd term over a common denominator and add. if you multiply through the top simplifies to 1